# Short research review: Applications of statistical physics investigating financial and other social systems

### 16 Mar 2024 v1

@@
A significant portion of our research focuses on modeling long-range memory phenomena and power-law statistics within complex systems. Notably, our approach diverges from traditional methodologies, as we employ Markovian models in place of non-Markovian alternatives. Despite this departure, we achieve similar statistical results by harnessing various nonlinear dependencies embedded within our models. In the context of SDEs, nonlinearity induces non-stationarity in the increments of the stochastic process, consequently leading to the emergence of spurious long-range memory [26, 27]. While our models differ from those representing true long-range memory, a crucial question emerges: Do our models accurately capture the observed memory in financial markets and potentially other socio-economic complex systems?
@@
The first component involves the development of a statistical test capable of distinguishing between spurious and genuine long-range memory. Our focus has centered on the Burst Duration Analysis (BDA) method [30, 9, 31, 32], demonstrating competitive performance relative to alternative approaches. The core tenet of this method posits that, for any one-dimensional Markovian random walk, the probability density function (PDF) of first-passage times should exhibit a power-law behavior with an exponent of $-3/2$," at least for specific durations. Deviations from this expected behavior may indicate the presence of authentic long-range memory. Nevertheless, it is worth noting that the effectiveness of this method may wane when dealing with non-one-dimensional stochastic processes, warranting further exploration, as exemplified in multidimensional scenarios [33].
The second component necessitates the selection of models that display both spurious and true long-range memory. Our prior research has introduced a diverse array of models reflecting spurious long-range memory. Consequently, our next steps involve formulating equivalent alternative models and exploring the characteristics of existing long-range memory models. In our publication [29], we delved into the estimation of long-range memory within fractional L’evy stable motion, instantiated through an ARFIMA(0,d,0) discrete process. This approach represents a generalization of fractional Brownian motion. However, additional models warrant consideration [34]. For instance, the multiplicative point process could be enhanced by replacing uncorrelated Gaussian noise with fractional Gaussian noise. Other correlation structures or variations in pulse durations could further extend the framework [35]. Alternative options encompass continuous-time random walk and complex contagion frameworks.

### 19 Mar 2024 v2

@@
A significant portion of our research focuses on modeling long-range memory phenomena and power-law statistics within complex systems. Notably, our approach diverges from traditional methodologies, as we employ Markovian models in place of non-Markovian alternatives. Despite this departure, we achieve similar statistical results by harnessing various nonlinear dependencies embedded within our models. In the context of SDEs, non-linearity induces non-stationarity in the increments of the stochastic process, consequently leading to the emergence of spurious long-range memory [26, 27]. While our models differ from those representing true long-range memory, a crucial question emerges: Do our models accurately capture the observed memory in financial markets and potentially other socio-economic complex systems?
@@
The first component involves the development of a statistical test capable of distinguishing between spurious and genuine long-range memory. Our focus has centered on the Burst Duration Analysis (BDA) method [30, 9, 31, 32], demonstrating competitive performance relative to alternative approaches. The core tenet of this method posits that, for any one-dimensional Markovian random walk, the probability density function (PDF) of first-passage times should exhibit a power-law behavior with an exponent of $-3/2$, at least for specific duration. Deviations from this expected behavior may indicate the presence of authentic long-range memory. Nevertheless, it is worth noting that the effectiveness of this method may wane when dealing with non-one-dimensional stochastic processes, warranting further exploration, as exemplified in multidimensional scenarios [33].
The second component necessitates the selection of models that display both spurious and true long-range memory. Our prior research has introduced a diverse array of models reflecting spurious long-range memory. Consequently, our next steps involve formulating equivalent alternative models and exploring the characteristics of existing long-range memory models. In our publication [29], we delved into the estimation of long-range memory within fractional Lèvy stable motion, instantiated through an ARFIMA(0,d,0) discrete process. This approach represents a generalization of fractional Brownian motion. However, additional models warrant consideration [34]. For instance, the multiplicative point process could be enhanced by replacing uncorrelated Gaussian noise with fractional Gaussian noise. Other correlation structures or variations in pulse duration could further extend the framework [35]. Alternative options encompass continuous-time random walk and complex contagion frameworks.